The ages of three sisters are consecutive odd integers. Seven more than the youngest sister's age is eight less than twice the oldest sister's age. How old is the middle sister?
Thank you for answering!

Found 2 solutions by amarjeeth123, Yetoo:
Answer by amarjeeth123(574)

Let the sisters ages be (x-2),x and (x+2)
where (x-2) is the younger sister's age
x is the middle sister's age
(x+2) is the elder sister's age
Seven more than the youngest sister's age is eight less than twice the oldest sister's age.
(x-2)+7=2(x+2)-8
x+5=2x-4
2x-x=5+4
x=9
The middle sister is 9 years old.
The younger sister is 7 years old.
The elder sister is 11 years old.

Answer by Yetoo(2)